Chapter 18, Problem 30 GO Suppose you want to determine the electric field in a certain region of space. You have a small object of known charge and an instrument that measures the magnitude and direction of the force exerted on the object by the electric field. (a) The object has a charge of +30.0 μC and the instrument indicates that the electric force exerted on it is 2SO μN, due east. What are the magnitude and direction of the electric field? (b) What are the magnitude and direction of the electric f eld if the object has a charge of-30.0 μC and the instrument indicates that the force is 10 N due west? E-
Answer #1

a) Charge on object, q = 30 micro C = 30*10^-6 C
   Force experienced, F = 25 micro N in east
   so let the electric field be E
   then qE = F
   E = F/q = 25/30 = 0.8333 N/C
   direction , along east direction
b) Charge on object, q = -30 micro C = -30*10^-6 C
   Force experienced, F = 10 micro N, in west
   so let the electric field be E
   then qE = F
   E = F/q = -10/30 = -0.3333 N/C
   so magnitude of E = 0.333 N/C
   direction , along east